OpenStudy (jchick):

Choose the equation of the line passing through the point (-3, -1) and perpendicular to y = one halfx + 3.

Nnesha (nnesha):

the equation you have is \[y = \frac{ 1 }{ 2 }x + 3\] you need perpendicular slope

OpenStudy (anonymous):

ERPENDICULAR LINES HAVE THE NEGATIVE RECIPROCAL SLOPES

Nnesha (nnesha):

do you know what is slope of that ^^^^ equation ??

OpenStudy (anonymous):

so your slope would be -2

OpenStudy (jchick):

No I have no idea what I am doing

OpenStudy (jchick):

Ok thanks

OpenStudy (anonymous):

now subtitute points: -1=-3*-2+b b=-1-6=-7

Nnesha (nnesha):

okay so slope intercept form is y =mx+b where m is slope which is always with x and b is y-intercept so what is m in this equation y =1/2x + b ??

OpenStudy (anonymous):

y=-2x-7 is final

OpenStudy (jchick):

a. y = -2x - 7 b. y = -2x +5 c. y = 2x + 5 d. y = 2x - 7

OpenStudy (anonymous):

well you see of course A

Nnesha (nnesha):

@dtuniyants can you pleases let her/him do this one at least give him/her chance

Nnesha (nnesha):

its against the coc hope you know it

OpenStudy (jchick):

No I just need an explanation

Nnesha (nnesha):

did you really help ?? this is not help its ENABLING

OpenStudy (jchick):

I do not like anyone's explanation better i just need help

OpenStudy (anonymous):

that right

OpenStudy (jchick):

I would like to understand @Nnesha

OpenStudy (jchick):

you think that i just want the answer no I am on the lesson and I am trying to figure it out

Nnesha (nnesha):

i really don't care about medals if you still want to learn it i would love to explain so you can understand

OpenStudy (jchick):

Yes please do

Nnesha (nnesha):

okay good so do you know what is slope of \[y = \frac{ 1 }{ 2 }x + 3\] ??

OpenStudy (jchick):

No i don't can you help me I don't know how to find it

Nnesha (nnesha):

slope intercept form is y =mx+b where m is slope which is always with x and b is y-intercept so what is m in this equation y =1/2x + b ??

OpenStudy (jchick):

1/2?

Nnesha (nnesha):

yes right

OpenStudy (jchick):

What do i do now?

Nnesha (nnesha):

so you need perpendicular slope perpendicular lines have negative reciprocal slopes so your slope is 1/2 you have to flop this fraction and change sign

Nnesha (nnesha):

flip**

OpenStudy (jchick):

-1/2

OpenStudy (jchick):

No sorry -2

Nnesha (nnesha):

yes that's right

Nnesha (nnesha):

now y =mx + b this is slope intercept form replace m by -2 replace x and y by their values and then solve for b